The range of a projectile is directly proportional to the square of its velocity. If a motorcyclist can make a jump of 247.0 feet by coming off a ramp at 71 mph, find the distance the motorcyclist could expect to jump if the speed coming off the ramp were increased to 76 mph. Round to the nearest tenth of a foot.
please help me solve! explain this to me please.
thanks
cz

Answer by nerdybill(7384)

The range of a projectile is directly proportional to the square of its velocity. If a motorcyclist can make a jump of 247.0 feet by coming off a ramp at 71 mph, find the distance the motorcyclist could expect to jump if the speed coming off the ramp were increased to 76 mph. Round to the nearest tenth of a foot.
.
When you see "directly proportiona" think
y = kx
where k is a constant
.
To find k, use:
"If a motorcyclist can make a jump of 247.0 feet by coming off a ramp at 71 mph"
247 = k(71)
247/71 = k
3.479 = k
.
Now, you have the realtion:
y = 3.479x
.
So, to answer:
"find the distance the motorcyclist could expect to jump if the speed coming off the ramp were increased to 76 mph"
y = 3.479x
y = 3.479(76)
y = 264.4 feet